A sum of 45000 invested will provide payments of 500 each at the end of every 2 months for 7 years find the under lying rate of interest compounded semi annualy


1
Expert's answer
2022-02-07T17:34:30-0500

Solution:

PV of payments "=500\times 3\times \dfrac{1-(\dfrac1{(1+\frac r2)^{7\times 2}})}{\frac r2}"

Invested sum = 45000

So, "45000=15000\times \dfrac{1-(\dfrac1{(1+\frac r2)^{7\times 2}})}{\frac r2}"

"\Rightarrow \dfrac {3r}2=1-\dfrac1{(1+\frac r2)^{14}}
\\ \Rightarrow \frac r2=0.326984
\\ \Rightarrow r=0.653968
\\\Rightarrow r=65.3968\%"